On a Physicians Licensure Exam with a population mean of 83 and standard deviation of 8. Jenny scores 4 points below the mean, Jeffrey had a score of x=84 and Jolly had a z-score of 0.30. Place these three students in order from the lowest to highest.


1
Expert's answer
2022-03-15T19:17:22-0400

"\mu=83, \ \ \sigma =8."

Jenny's score: "x_1=\mu-4=83-4=79."

Jeffrrey's score: "x_2=84."

Jolly's z-score: "z_3=\frac{x_3-\mu}{\sigma},"

Jolly's score:

"x_3=z_3\sigma+\mu=0.3\cdot8+83=85.4."

The students in order from the lowest to highest: Jenny, Jeffrrey, Jolly.